OpenStudy (anonymous):

The displacement of a body with time t is related as S=ae^6t - be^-4t. Find the rate of change of displacement with respect to time t.

OpenStudy (sidsiddhartha):

rate of change of displacement with respect to time t. means u have to find \[\frac{ ds }{ dt }\]

OpenStudy (anonymous):

how to ds/dt S=ae^6t - be^-4t ?

OpenStudy (sidsiddhartha):

s is given now just differentiate it with respect to t

geerky42 (geerky42):

You just need to apply chain rule for exponential function. do you know that rule?

OpenStudy (sidsiddhartha):

and (ds/dt) means the velocity of the particle

OpenStudy (anonymous):

\[S=ae ^{6t}-be ^{-4t}\]

geerky42 (geerky42):

\(\Large \dfrac{d}{dt}e^t = e^t\)

OpenStudy (sidsiddhartha):

hmm then u can find (ds/dt) can't u :)

geerky42 (geerky42):

Have you been taught calculus before?

OpenStudy (anonymous):

Yup

OpenStudy (sidsiddhartha):

\[\huge \frac{ d }{ dt}e^{at}=ae^{at}\] as @geerky42 said

OpenStudy (anonymous):

so what now ..!

OpenStudy (sidsiddhartha):

now just simple differentiation using chain rule \[\large \frac{ ds }{ dt }=\frac{ d }{ dt }[ae^{6t}-be^{-4t}]\]

OpenStudy (anonymous):

\[ S \prime= ae ^{6t}.(6)-be ^{-4t}.(-4)\] right?

OpenStudy (sidsiddhartha):

yup good :)

OpenStudy (anonymous):

what's next?

OpenStudy (sidsiddhartha):

nothing ds/dt is evaluated :)

OpenStudy (anonymous):

But the question says find the rate of change of displacement with respect to time t. :|

OpenStudy (sidsiddhartha):

yeah let s is the displacement then rate of change of displacement with respect to time t. is \[\frac{ \Delta s }{ \Delta t }=\frac{ ds }{ dt}\] ok???

OpenStudy (anonymous):

ammmmmm no I don't understand

OpenStudy (anonymous):

I will come back BRB

OpenStudy (sidsiddhartha):

let displacement of a particle at time T1 is S1 and sometime later at T2 it is S2 then change in time =(T2-T1) and change in displacement=(S2-S1) so rate of change of displacement with respect to time is \[\frac{ S2-S1 }{ T2-T1}=\frac{ \Delta S }{ \Delta T}=ds/dt\] do u get it now?

OpenStudy (kainui):

I don't think soso realizes that he has the answer already, but I'm not sure what he doesn't understand.

OpenStudy (sidsiddhartha):

yeah mee too

OpenStudy (anonymous):

:(